A 1200 kg car traveling at 15 m/s due east collides with a 720 kg car traveling at 21 m/s due north. As a result of the collision the two cars lock together what is the velocity immediately after the collision?


1
Expert's answer
2021-05-10T09:18:55-0400


 conservation of momentum for inelastic collision tell us

Initial momentum = final momentum

mu+MU=(m+M)V

1200×15i+720×21j=(1200+720)V

V=(18000i+720x21j)/1920

V=9.375i+7.875j

​Final velocity when they stick together is 9.375i+7.875j

It is perpendicular to both each and north direction

Resultant velocity is =√(9.375^2+7.875^2)

=√(87.89+62.01)

=12.24m/s

​